OpenStudy (truwhovian11):

WILL MEDAL AND FAN FIRST PERSON TO HELP ME 1 OF 5 QUESTIONS Multiply and simplify. 6 1/4 x 4/5 A.1/5 B.5 C.1/4 D.4

OpenStudy (dayakar):

6 1/4 *4/5 =( 6*4 +1)/4 *4/5 = 25/4 * 4/5 can u do remaining part

OpenStudy (truwhovian11):

B.5 ?

OpenStudy (dayakar):

yes ,ur correct

OpenStudy (truwhovian11):

thank you i have more questions tho

OpenStudy (dayakar):

write quickly!

OpenStudy (truwhovian11):

ok